### v2

@@ 2 Results
As shown in Table 1, genera imbalance strongly influenced richness estimation error ($RMSE_{\alpha}$), though not uniformly across cities. In highly uneven forests such as Washington and Seattle, $RMSE_{\alpha}$ exceeded within-city variability, underscoring the difficulty of estimating richness under strong dominance patterns. By contrast, in more balanced systems such as Los Angeles, Denver, and New York, errors were smaller than or comparable to observed standard deviations, and in several cases outperformed a naïve mean-predictor baseline (see Appendix D). Despite these differences, the V-score remained consistent across all datasets (0.397–0.498), indicating that clustering preserved the overall partitioning structure even when richness estimates deviated. These findings suggest that richness is not the primary strength of our approach, but performance is competitive in balanced forests and improves over statistical baselines.
@@
In our datasets, Simpson diversity is bounded between 0 and 1, while Shannon diversity can reach values of about 3.5 (see Section 4.1). To enable direct comparison across indices and cities, Shannon values were normalized by their relative maximum, placing both metrics on a common [0,1] scale. In this context, Wasserstein distances below 0.05 indicate very strong alignment with ground truth distributions, while values near 0.1 represent moderate but still ecologically meaningful agreement. Cities such as Washington (0.133 for Shannon, 0.033 for Simpson) and San Francisco (0.034 for Shannon, 0.028 for Simpson) demonstrated strong alignment for both indices, with proportionally smaller errors for Simpson. Calgary, by contrast, exhibited the weakest alignment across metrics. We hypothesize that this may be linked to more aggressive visual obfuscation apparent in Canadian Street View imagery compared to the United States, where stronger privacy safeguards and stricter blurring protocols are enforced [29]. Fig. 1 illustrates this by comparing predicted and observed values in Los Angeles (Shannon) and New York (Simpson). A full set of city-level results, including baseline comparisons and variability relative to observed standard deviations, is provided in Appendix D, which contextualizes how performance varies across urban forest structures.
@@
One potential concern with unsupervised clustering is that lower distributional errors may arise not from genuine recovery of community structure but from artificial entropy inflation, e.g., through oversplitting or the proliferation of singleton clusters. To assess this, we performed a dedicated diagnostic analysis (Appendix C) that examined oversplitting factors, purity, evenness shifts, and rank–abundance correlations between predicted and observed communities. Across all cities, the singleton ratio was zero, oversplitting factors remained close to one, and evenness shifts were modest, with New York even showing a slight decrease. At the same time, rank–abundance correlations were consistently high ($\rho\approx 0.9$), indicating that predicted clusters preserved the overall shape of community distributions. These findings confirm that the improved alignment of Shannon and Simpson diversity under the unsupervised approach is not an artifact of inflated entropy, but rather reflects ecologically meaningful recovery of biodiversity structure.
Spatial fidelity results.
City Richness Shannon Simpson
Calgary 0.549 0.617 0.612
New York 0.713 0.766 0.769
Columbus 0.577 0.652 0.633
Denver 0.823 0.860 0.836
Los Angeles 0.524 0.566 0.519
Seattle 0.683 0.783 0.764
Washington 0.726 0.823 0.816
San Francisco 0.630 0.761 0.762
Average 0.653 0.729 0.714
@@
Overall, the method achieved low Wasserstein distances for both Shannon and Simpson diversity across diverse urban forest conditions, including in highly imbalanced contexts. High Simpson accuracy suggests the framework effectively captured dominant genera distributions, while strong Shannon performance indicates robustness to abundance evenness patterns. Taken together, these results demonstrate that our approach can accurately recover both abundance distributions and spatial diversity patterns from entirely unlabeled street-level imagery, directly addressing the lack of scalable, label-free biodiversity mapping methods highlighted in the Introduction.
@@ 4.1 Data Description
Collectively, these differences position the dataset used as a diverse and challenging benchmark to test the generalizability of unsupervised clustering methods in urban biodiversity assessment. A detailed pseudo-code of the pipeline is provided in Appendix A (Algorithm 1), and the complete implementation, including hyperparameter tuning and preprocessing scripts, is available on Github.
Figure 3: Generation of pseudo-clusters using spatial priors. Street-level images are sampled within local grids, visual embeddings are extracted, and spatial proximity guides grouping into initial clusters.
@@ 4.2.1 Pseudo Clustering
As an initial step, we group TaxaBind spatial embeddings into pseudo-clusters using the HDBSCAN algorithm [40], leveraging the tendency of urban trees to be planted in structured, often linear arrangements. This approach, illustrated in Fig. 3, typically results in taxonomically pure clusters due to the spatial regularity introduced by common urban planting strategies such as row planting along streets or uniform spacing in public parks.
@@ 4.2.3 Re-grouping and Re-assignment
The re-grouping and re-assignment process (Fig. 5) targets samples eliminated during the outlier detection stage, which were initially clustered based primarily on spatial proximity rather than visual coherence. Since such pseudo-clusters may group together visually dissimilar trees due to their physical arrangement, the eliminated members are revisited to assess whether they form more semantically consistent groupings.
@@
Figure 5: Outliers are either re-grouped into new visually coherent clusters or reassigned to existing ones, using tuned similarity thresholds.
@@ Appendix B Ablation analysis
Appendix C Supervised Diagnostics
@@
Appendix D Comparisons with Statistical Baselines
@@
Appendix E Validation of extracted visual features

### v3

@@ 2 Results
As shown in Table 1, genera imbalance strongly influenced richness estimation error ($RMSE_{\alpha}$), though not uniformly across cities. In highly uneven forests such as Washington and Seattle, $RMSE_{\alpha}$ exceeded within-city variability, underscoring the difficulty of estimating richness under strong dominance patterns. By contrast, in more balanced systems such as Los Angeles, Denver, and New York, errors were smaller than or comparable to observed standard deviations, and in several cases outperformed a naïve mean-predictor baseline (see Appendix E). Despite these differences, the V-score remained consistent across all datasets (0.397–0.498), indicating that clustering preserved the overall partitioning structure even when richness estimates deviated. These findings suggest that richness is not the primary strength of our approach, but performance is competitive in balanced forests and improves over statistical baselines.
@@
In our datasets, Simpson diversity is bounded between 0 and 1, while Shannon diversity can reach values of about 3.5 (see Section 4.1). To enable direct comparison across indices and cities, Shannon values were normalized by their relative maximum, placing both metrics on a common [0,1] scale. In this context, Wasserstein distances below 0.05 indicate very strong alignment with ground truth distributions, while values near 0.1 represent moderate but still ecologically meaningful agreement [29]. Cities such as Washington (0.133 for Shannon, 0.033 for Simpson) and San Francisco (0.034 for Shannon, 0.028 for Simpson) demonstrated strong alignment for both indices, with proportionally smaller errors for Simpson. Calgary, by contrast, exhibited the weakest alignment across metrics. We hypothesize that this may be linked to more aggressive visual obfuscation apparent in Canadian Street View imagery compared to the United States, where stronger privacy safeguards and stricter blurring protocols are enforced [30]. Fig. 1 illustrates this by comparing predicted and observed values in Los Angeles (Shannon) and New York (Simpson). A full set of city-level results, including baseline comparisons and variability relative to observed standard deviations, is provided in Appendix E, which contextualizes how performance varies across urban forest structures.
@@
One potential concern with unsupervised clustering is that lower distributional errors may arise not from genuine recovery of community structure but from artificial entropy inflation, e.g., through oversplitting or the proliferation of singleton clusters. To assess this, we performed a dedicated diagnostic analysis (Appendix D) that examined oversplitting factors, purity, evenness shifts, and rank–abundance correlations between predicted and observed communities against a supervised model using linear probing on BioCLIP [23] model. Across all cities, the singleton ratio was zero, oversplitting factors remained close to one, and evenness shifts were modest, with New York even showing a slight decrease. At the same time, rank–abundance correlations were consistently high ($\rho\approx 0.9$), indicating that predicted clusters preserved the overall shape of community distributions. These findings confirm that the improved alignment of Shannon and Simpson diversity under the unsupervised approach is not an artifact of inflated entropy, but rather reflects ecologically meaningful recovery of biodiversity structure.
@@
Overall, the method achieved low Wasserstein distances for both Shannon and Simpson diversity across diverse urban forest conditions, including in highly imbalanced contexts. High Simpson accuracy suggests the framework effectively captured dominant genera distributions, while strong Shannon performance indicates robustness to abundance evenness patterns. Taken together, these results demonstrate that our approach can accurately recover both abundance distributions and spatial diversity patterns with strong spatial fidelity (see Appendix C) from entirely unlabeled street-level imagery, directly addressing the lack of scalable, label-free biodiversity mapping methods highlighted in the Introduction.
@@ 4.1 Data Description
Collectively, these differences position the dataset used as a diverse and challenging benchmark to test the generalizability of unsupervised clustering methods in urban biodiversity assessment. A detailed pseudo-code of the pipeline is provided in Appendix B (Algorithm 1), and the complete implementation, including hyperparameter tuning and preprocessing scripts, is available on Github.
@@ 4.2.1 Pseudo Clustering
As an initial step, we group TaxaBind spatial embeddings into pseudo-clusters using the HDBSCAN algorithm [41], leveraging the tendency of urban trees to be planted in structured, often linear arrangements. This approach typically results in taxonomically pure clusters due to the spatial regularity introduced by common urban planting strategies such as row planting along streets or uniform spacing in public parks.
@@ 4.2.3 Re-grouping and Re-assignment
The re-grouping and re-assignment process targets samples eliminated during the outlier detection stage, which were initially clustered based primarily on spatial proximity rather than visual coherence. Since such pseudo-clusters may group together visually dissimilar trees due to their physical arrangement, the eliminated members are revisited to assess whether they form more semantically consistent groupings.
@@ Appendix B Ablation analysis
Appendix C Spatial fidelity results.
City Richness Shannon Simpson
Calgary 0.549 0.617 0.612
New York 0.713 0.766 0.769
Columbus 0.577 0.652 0.633
Denver 0.823 0.860 0.836
Los Angeles 0.524 0.566 0.519
Seattle 0.683 0.783 0.764
Washington 0.726 0.823 0.816
San Francisco 0.630 0.761 0.762
Average 0.653 0.729 0.714
Appendix D Supervised Diagnostics
@@
Appendix E Comparisons with Statistical Baselines
@@
Appendix F Validation of extracted visual features